Stephanie Corporation sells a single product. Budgeted sales for the year are anticipated to be 688,000 units, estimated beginning inventory is 110,000 units, and desired ending inventory is 87,000 units. The quantities of direct materials expected to be used for each unit of finished product are given below. Material A 0.50 lb. per unit @ $0.74 per pound Material B 1.00 lb. per unit @ $1.91 per pound Material C 1.20 lb. per unit @ $0.88 per pound The dollar amount of material A used in production during the year is a. $702,240 b. $1,270,150 c. $246,050 d. $254,560
Answer #1

Production units = 688000+87000-110000 = 665000 Units

Dollar amount of material A used in production = 665000*.50*0.74 = 246050

So answer is c) $246050